Pls I need help with this also
Specify the domain and range of the relation
x^2/18+(y-4)^2/4=1/2

Answer by venugopalramana(3286)

Pls I need help with this also
Specify the domain and range of the relation
x^2/18+(y-4)^2/4=1/2
(y-4)^2 = 4[0.5-x^2/18]=(4/18)[9-x^2]
since (y-4)^2 is always non -ve,9-x^2 shall be >=0
9>=x^2
3>=|x|
|x|<=3....is the domain..that is (-3,3)is the domain.
the function has a minimum at x=3....then y-4=0....y=4
hence range is [4,infinity)